# Privacy Preservation Techniques (PPTs) in IoT Systems: A Scoping Review and Future Directions Author’s draft for soliciting feedback - February 20, 2025

### v1

@@ VIII-B Technology used in building PETs
PETs are tools used to create PPTs in a system. These PETs are built using innovative technologies such as blockchain, machine learning/deep learning, and post-quantum algorithms, to mention a few, as shown in Figure 9. Some of these technologies are explained as follows:
VIII-B1 Blockchain
@@
Blockchain technology decentralizes the control system, fostering a distributed and peer-to-peer transaction environment among anonymous parties within a digital ledger system [300] [301]. It enhances privacy in IoT systems through smart contracts, which enable trust transactions that prevent identity disclosure and tracking [302].
Blockchain has emerged as a novel protective layer for the internet and its applications. Privacy is achieved through the provision of authentication, validation, and tracing mechanisms, serving as robust safeguards for assets within the IoT system. [300].
VIII-B2 Cryptography
@@
This approach encodes the data to be accessed by only authorized users. An example is Homomorphic protection, which allows encrypted data to be computed at its encrypted state without decrypting it [229] [234].
VIII-B3 Machine and Deep Learning (MD/DL)
@@
The introduction of AI has allowed the implementation of ML/DL in IoT systems for privacy protection. However, their implementation circle around protecting data in process at the nodes or the cloud (service provider end), as seen in federated learning(FL) applications in IoT privacy protection [298, 290, 284, 235, 212, 140, 23, 118, 87, 151, 27, 24].
VIII-B4 Mathematical Model
@@
Some studies proposed novel mathematical models to protect privacy in the IoT environment. Most of these models center on protecting data in transit during communication among IoT system components. The proposed mathematical model either provides cryptography or access control to the system [32, 270, 187, 97, 100, 220].
VIII-B5 Optimization
@@
VIII-B6 Post-Quantum Algorithm
This approach utilizes a post-quantum algorithm to ensure quantum-safe privacy protection in IoT. Liu et al. [133] proposed a lattice-based encryption scheme to ensure the encrypted data are quantum-safe. This is the only paper in this study that proposed post-quantum encryption for privacy protection.
VIII-B7 Infrastructure-based
@@
VIII-B8 Informed Consent Management
It is the the only non-technological approach in our study. It relies on policies and laws governing data privacy, such as GDPR, Personal Information Protection and Electronic Documents Act (PIPEDA), California Consumer Privacy Act (CCPA), and Brazilian General Data Protection Act (LGPD), as outlined by regulatory bodies [278, 271, 295].
Our observations: Some studies combine some of these PET building technologies to achieve multiple privacy preservation goals. Examples are Singh et al. [140], Fang et al. [118], and Zhao et al. [151] all use ML and blockchain to provide privacy preservation in IoT. Also, Salim et al. [27], and Hindistan [25] both utilize ML and DP to build PET in IoT systems.
VIII-C IoT Applications and PPTs
The literature typically focuses on common IoT applications in transportation, home/building, city, health, and industry. However, IoT applications are expanding into social networks and multimedia, as seen in social IoT and the Internet of Multimedia Things (IoMT).
VIII-C1 Social IoT (S-IoT)
This can be seen in the mutually coordinated social cooperation that exists among IoT devices, with their connection strength depending on interaction levels. This connections are built on trust and privacy [219][186] [218]. These interactions can leverage platforms like Web 3.0, as seen in social media applications such as SM-IoT [27] [296].
VIII-C2 Internet of Multimedia of Things (IoMT)
@@
Figure 6: Papers on PPT implementation in IoT applications and Technology use to build PETs from 2010 to the first quarter of 2023
IoT Application Edge Fog Cloud Local Edge and Cloud Fog and Cloud Fog, Edge and Cloud Local, Edge, Fog and Cloud Local and Cloud Not Mentioned
Smart Home [151], [264], [265], [131], [83], [72] [270], [148] [278],[236], [223], [224],[288], [71] - [66] - - - - [269], [246], [105], [157], [276], [115]
Industrial IoT (IIoT) [212] [26] [252], [229], [25] [133],[yin2017location] [70] - [156] - - [251], [87], [211]
Smart Grid [165],[215] - [158], [237], [203] - [213] - - - - [289], [92], [160]
Medical IoT [234],[146], [79], [147],[74], [149], [78] [94], [238],[73] [116], [140], [84], [240], [245], [111], [68],[129],[235],[166],[101],[195],[258],[259],[303],[239],[117],[249] [191] [80] - - - - [142],[170],[214],[114]
Social IoT (S-IoT) [219] - [186], [183] - - - - - - [296],[27],[218]
IoT of Vehicle (IoV) [222] [185],[242],[89],[176],[199] [90],[206],[253],[188],[280],[273],[260],[107] [180] [241], [81], [267] [75] - - - [28],[189],[201], [192],[102],[291],[193],[150],[263]
Wearable IoT - - [190], [293], [178] [196] - - - - - [97],[298]
Smart City [113],[272],[210] [204], [182] [32],[86],[283] - [217] - - - [173],[254] [208], [152],[187]
Internet of Multimedia of things (IoMT) - [285] [123] - - - - - - [297]
IoT in Agriculture - - [125] - - - - - - -
IoT in Education - - [167] - - - - - - -
IoT in Ecommerce - - - - - - - - - [99]
General IoT [169],[91],[230],[279],[118],[244],[110],[163],[154],[209],[247],[177],[139],[198], [216], [180], [181], [184], [82], [299], [76], [287], [132], [77] Fog [205],[243],[153], [161],[164],[295],[23],[281], [286], [277] [141],[250],[88],[268],[119],[126],[202],[badsha2018designing],[162],[112],[98],[225],[256],[232],[233],[103],[292],[220],[221],[109] [207],[290],[271],[69], [231], [282] [93],[284],[127] [257] - [24] [227] [85],[122],[143][22],[134],[226],[95],[96], [135],[255],[274],[136],[130],[137],[175],[120],[228],[144],[155],[275],[100],[104],[106],[128],[124],[145],[294],[168],[138],[197],[171],[172],[67],[200],[108],[261],[262],[159],[121], [266], [179], [77]
Table IV: Paper on PPTs and Layer of Implementation in IoT architecture from 2010 to the first quarter of 2023.
@@
User Data Privacy 164 [32] [92] [143] [22] [113] [119] [151] [237] [160] [187] [280] [227] [269] [189] [94] [243] [201] [126] [153] [202] [223] [264] [190] [161] [114] [245] [271] [96] [135] [246] [255] [162] [123] [110] [130] [174] [137] [111] [224] [191] [98] [209] [120] [288] [238] [283] [257] [99] [144] [273] [204] [192] [275] [100] [101] [102] [291] [284] [105] [66] [193] [167] [194] [157] [146] [258] [145] [259] [68] [129] [293] [294] [276] [260] [168] [25] [107] [220] [138] [197] [235] [221] [170] [296] [211] [242] [198] [169] [199] [152] [159] [73] [118] [297] [27] [172] [213] [125] [212] [67] [239] [298] [262] [109] [214] [69] [263] [166] [156] [88] [268] [230] [24] [279] [118] [122] [278] [26] [156] [70] [158] [116] [140] [84] [240] [142] [117] [219] [241] [90] [272] [205] [250] [85] [141] [121] [265] [71] [215] [72] [216] [181] [184] [131] [178] [79] [147] [80] [182] [81] [183] [148] [267] [82] [282] [249] [299] [76] [179] [149] [83] [132] [77] [78]
Device Data Privacy 13 [290] [134] [93] [274] [164] [228] [106] [171] [91] [207] [115] [133] [87]
User and Device Data Privacy 24 [208] [173] [163] [97] [112] [233] [103] [104] [295] [139] [23] [281] [270] [200] [108] [261] [236] [251] [252] [289] [234] [186] [74] [203]
Location Privacy 20 [226] [188] [95] [175] [225] [176] [195] [124] [177] [196] [285] [155] [150] [218] [28] [89] [180] [231]
General Privacy 22 [244] [136] [154] [165] [256] [232] [210] [127] [292] [128] [247] [229] [222] [185] [206] [253] [86] [217]
Table V: Publications on privacy protection types in IoT systems from 2010 to the first quarter of 2023.
Our observations: According to Figure 6, the number of papers on PET building technology specific to IoT applications in agriculture, education, and e-commerce is the lowest. In contrast, the fields of Smart Homes, Medical IoT, and Industrial IoT (IIoT) contain a significant number of papers within our study, as illustrated in Figure 6. Additionally, some papers propose PET building technologies that could be applied across all IoT applications. However, this general approach may not be suitable for specific IoT applications, given the differences in architectures, protocols, resource requirements, and communication processes, as noted by Sana et al. [304].
VIII-D PPT implementation in IoT Application and System Architecture
PPT implementation in IoT has been proposed across different IoT architectures. These architectures have different layers where PPTs can be deployed. This study focus on the implementation of PPTs in IoT computing layers as shown in Figure 7.
PPT can be locally deployed within the IoT system at the aggregate (gateway) or at the edge close to the IoT system (which is common in Fifth-Generation (5G) enabled IoT systems). Fog and cloud-based architecture are also commonly used today, and most privacy studies have used some of these layers in IoT architecture to deploy PPTs.
Our observations: Some proposed studies on PPTs are deployed across two or more layers of the IoT architectures. Table IV shows the categorization of papers on PPTs and IoT computing layers in which they are deployed. Research work on end-to-end PPT deployment in IoT architecture that cuts across local, fog, edge, and cloud is limited, as shown in Table IV. Also, it was observed that the cloud is the most used platform where PPTs have been deployed in IoT architecture, followed by edge, then fog. However, many papers did not mention which computing layer they deployed PPT in their publications, as shown in Figure 7.
Figure 7: Publications on PPTs implementation at different computing layers in IoT architecture from 2010 to the first quarter of 2023
VIII-E Privacy Types provided by PPTs in IoT systems
As gathered from this study, the three main privacy types focused on in most papers are user privacy, device privacy, and location privacy. User data privacy protects sensitive data such as Personally Identifiable Information (PII), Protected Health Information (PHI), or any sensitive information that can reveal the identity of a user. The protection of data stored, processed, or transmitted across devices that could reveal either user or device identity in IoT is called device data privacy. Location privacy protects IoT user protection against privacy threats like localization and tracking. This threat also applies to mobile IoT devices like wearable or stationed devices.
Apart from these three privacy types common in IoT systems, some studies have proposed PPTs that simultaneously mitigate user and device privacy. Meanwhile, some studies have presented the general privacy of the entire IoT system without any specific privacy type.
Our observations: This study presents different privacy types proposed from the literature on PPT implementation in IoT systems. User data privacy has the highest number of papers focusing on user data protection. Few papers proposed PPT to protect device data and location privacy, as shown in Table V and Figure 8. This answers our research question RQ6.
Our study observed that device data privacy proposed by most studies still focuses on user data being stored, processed, or transmitted across IoT devices. This data could comprise both user and device data. Therefore, to protect device identity, sensitive device data such as IP address, MAC address, Operating system, and serial number, to mention a few that could reveal device identity, need to be protected. These sensitive data associated with IoT devices could be linked with other data to identify the device in the IoT system.
Figure 8: Papers on Privacy protection types in IoT system from 2010 to the first quarter of 2023
IX Step 5: Collating, Summarizing, and Reporting Results that Address Research Questions
Using the PRISMA-ScR guidelines for selecting relevant papers, our analysis reveals a steady increase in publications on PPTs from 2013 to 2019, with 2019 having the highest number of publications, as shown in Figure 4. There was a noticeable decline in 2020, probably due to the global pandemic. However, publication rates rebounded and continued to rise from 2021 through the first quarter of 2023, when the paper collection was concluded. The following subsections present results that address the research questions outlined in section V.
IX-A RQ1: What technologies are prominently used to develop
PETs in IoT systems? To address RQ1, this study analyzed seven technologies used to build Privacy-Enhancing Technologies (PETs)—Blockchain, Cryptography, Machine and Deep Learning, Mathematical Models, Optimization, Infrastructure-Based techniques, and post-quantum algorithms. Informed consent management is the only non-technical approach to privacy preservation, as gathered from our study. It has also been revealed that cryptography is the most prevalent technology used for privacy preservation in IoT applications, followed closely by blockchain. As observed in this study, the post-quantum algorithm is primarily applied to Industrial IoT (IIoT), indicating its pioneering role and potential for future research in IIoT privacy.
IX-B RQ2: What are the dominant PPT’s privacy goals in IoT applications?
Our study’s research question (RQ2) investigates the dominant goals of PPTs across IoT applications. Our analysis identifies various PPT goals, including anonymity, data flow privacy, access control, privacy awareness, data isolation, and data minimization, as shown in Table III. Among these goals, anonymity, particularly through methods like data masking and data perturbation, emerges as the most prevalent goal in the literature, with a significant percentage of publications focusing on these techniques.
IX-C RQ3: What are the commonly known and recent IoT applications added to the literature?
IoT applications commonly span cities, transportation, health, and agriculture. However, new application areas like Social IoT(S-IoT) and Internet Multimedia of Things (IoMT) are gaining traction as shown in Figure 6. These new platforms directly interact with users and are particularly vulnerable to privacy threats due to their extensive internet exposure.
IX-D RQ4: What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT most implemented from the literature?
IoT applications have increased recently, with medical IoT having the highest PPTs implementation as seen in Figure 6. This shows the importance of privacy in medical IoT systems. This level of importance is a result of sensitive medical data transactions involved in the MIoT system containing data such as Personally Identifiable Information (PII) and Protected Health Information (PHI). In case of a data breach, these sensitive medical data can reveal the identity of an individual in relation to their medical condition. Furthermore, our observation shows that PPTs are more implemented around the cloud layer in the IoT architecture in most IoT applications compared to other computing layers, as shown in Table IV.
@@ IX-C RQ3: What are the commonly known and recent IoT applications added to the literature?
IX-E RQ5: What privacy types are the main focus of most PPTs in IoT systems?

### v2

@@ 8.2 Technology used in building PETs
PETs are tools used to create PPTs in a system. These PETs are built using innovative technologies, such as blockchain, machine learning/deep learning, and post-quantum algorithms, among others, as shown in Figure 4. Some of these technologies are explained as follows:
8.2.1 Blockchain
@@
Blockchain technology decentralizes the control system, fostering a distributed and peer-to-peer transaction environment among anonymous parties within a digital ledger system [280] [60]. It enhances privacy in IoT systems through smart contracts, which enable trust transactions that prevent identity disclosure and tracking [208]. Blockchain has emerged as a novel protective layer for the internet and its applications, achieving privacy through the provision of authentication, validation, and tracing mechanisms, which serve as robust safeguards for assets within the IoT system. [280].
8.2.2 Cryptography
@@
This approach encodes the data to be accessed by only authorized users. An example is Homomorphic encryption, which allows encrypted data to be computed at its encrypted state without decrypting it [267] [511].
8.2.3 Machine and Deep Learning (MD/DL)
@@
Both ML and DL are subsets of AI; however, ML enables computers to learn and adapt patterns from given data with little human intervention [62]. DL is more efficient than ML, as it utilizes artificial multi-layers within an artificial neural network to learn and adapt patterns similar to the human brain [430]. The introduction of AI has allowed the implementation of ML/DL in IoT systems for privacy protection. However, their implementation circles around protecting data in process at the nodes or the cloud (service provider end), as seen in federated learning(FL) applications in IoT privacy protection [79, 626, 247, 201, 549, 472, 439, 150, 185, 628, 422, 87].
8.2.4 Mathematical Model
@@
Some studies have proposed novel mathematical models to protect privacy in the IoT environment. Most of these models center on protecting data in transit during communication among IoT system components. The proposed mathematical model either provides cryptography or access control to the system [170, 44, 489, 623, 412, 589].
8.2.5 Optimization
@@
8.2.6 Post-Quantum Algorithm
This technology is used to enhance privacy in IoT systems by hardening defenses against quantum-based privacy breaches. There are a few publications on the post-quantum use of IoT systems; we identified two of the most relevant papers in this study. The first paper of 2023 by Liu et al. [287] employs a post-quantum algorithm in industrial IoT (IIoT) to ensure quantum-safe privacy protection. They proposed a lattice-based encryption scheme to ensure the encrypted data is quantum-safe. The second paper by Yi et al. [594] uses post-quantum cryptography as the security foundation for an IoT-based cryptocurrency (an example of an IoT application in e-commerce). The focus is on enabling IoT devices to participate in cryptocurrency markets and transact securely. This preserves the user’s privacy using quantum-resistant encryption algorithms against both classical and quantum adversaries [594].
8.2.7 Infrastructure-based
@@
8.2.8 Informed Consent Management
This is the only non-technological approach in our study. It relies on policies and laws governing user data privacy, with focus on consent management and user control in alignment with privacy regulations such as GDPR, Canadian Personal Information Protection and Electronic Documents Act (PIPEDA), California Consumer Privacy Act (CCPA), and Brazilian General Data Protection Act (LGPD), as outlined by regulatory [372, 86, 19, 410, 372]. However, other present regulatory frameworks that emphasize user control and informed consent that protect privacy rights, as seen in [376].
Our observations: Some studies combine these PET building technologies to achieve multiple privacy-preserving goals. Examples are Singh et al. [472], Fang et al. [150], and Zhao et al. [628], all of which use ML and blockchain to provide privacy preservation in IoT. Additionally, Salim et al. [422] and Hindistan [194] both employ ML and DP to develop PET in IoT systems. Moreover, some authors combine three technologies to achieve privacy preservation in IoT systems. For example, blockchain, ML/DL (especially federated learning), and Cryptography were employed by Khan et al. [242], Guduri et al. [179], and Cheng et al. [105]. Others, such as Rehman et al. [407], employ federated learning, cryptography, and an optimization approach. In contrast, Jagdale et al. [216] employ a blockchain, a convolutional neural network, and a post-quantum algorithm to preserve privacy in IoT systems.
8.3 IoT Applications and PPTs
The literature typically focuses on common IoT applications in transportation, home/building, city, health, and industry. However, IoT applications are expanding into social networks and multimedia, as seen in social IoT, the Internet of Multimedia Things (IoMT), and Artificial Intelligence of Things.
8.3.1 Social IoT (S-IoT)
This can be seen in the mutually coordinated social cooperation that exists among IoT devices, with their connection strength depending on interaction levels. These connections are built on trust and privacy [449][21] [205]. These interactions can leverage platforms like Web 3.0, as seen in social media applications such as SM-IoT [422] [423].
8.3.2 Internet of Multimedia of Things (IoMT)
@@
8.3.3 Artificial Intelligence of Things (AIoT)
The AIoT combines Artificial Intelligence (AI) and IoT in which the Smart device, not only collects information, but understands it, learn this information, and makes decisions solely without human intervention through the help of AI [545, 528]. AI in IoT is typically implemented at edge devices, where raw data is processed locally and differential privacy is applied, adding controlled noise to each device’s model before sharing. The device shares the learned data with the central location while leaving the raw data on the device [545, 528]. Park et al. [368] employ the lightweight key exchange and hashing to achieve strong encryption without exposing sensitive data. In their approach to privacy preservation in AIoT, the authors Ivanovska et al. [213] introduce a comprehensive privacy-by-design approach to implement a PPT directly on the hardware and system architecture preserve the privacy of the data collected by these hardware and an AI processing pipeline [213]. The application of AIoT has cut across IoT applications such as Smart Homes [528], Smart City [213], and MIoT [368].
Our observations: According to Figure 12, the number of papers on PET building technology specific to IoT applications in agriculture, education, and e-commerce is the lowest. In contrast, the fields of Smart Homes, Medical IoT, and Industrial IoT (IIoT) contain a significant number of papers within our study, as illustrated in Figure 12. Additionally, some papers propose PET building technologies that could be applied across all IoT applications. However, this general approach may not be suitable for specific IoT applications, given the differences in architectures, protocols, resource requirements, and communication processes, as noted by Sana et al. [209]. Moreover, the introduction of AIoT in IoT environment, gives brain to the IoT devices to learn, reason and make decision without human intervention. However, this may open the door to another aspect of privacy threat that could impact the decisions of these IoT devices.
8.4 PPT implementation in IoT Application and System Architecture
PPT implementation in IoT has been proposed across different IoT architectures. These architectures have different layers where PPTs can be deployed. We focus on the implementation of PPTs in IoT computing layers as shown in Figure 3.
PPT can be locally deployed within the IoT system at the aggregate (gateway) or at the edge device. Fog and cloud-based architectures are also commonly used today, and most privacy studies have utilized some of these computing layers in IoT architectures to deploy PPTs. There are many vulnerabilities associated with specific computing layers, as some of these layers are managed by third parties, which could result in jurisdictional risks privacy threats [359]. Additionally, some cloud computing platforms hosted outside the privacy regulations applicable to the location of the IoT system’s deployment may not comply with those regulations. This is one of the privacy threats that denies user control over their data, as mentioned in Canadian PIPEDA fair information principle [361].
Our observations: Some proposed studies on PPTs are deployed across two or more layers of the IoT architectures. Table LABEL:table:IoT-Architecturetable shows the categorization of papers on PPTs and IoT computing layers in which they are deployed. Research on end-to-end PPT deployment in IoT architectures, which span local, fog, edge, and cloud, is limited, as shown in Table LABEL:table:IoT-Architecturetable. Additionally, it was observed that the cloud is the most widely used platform for deploying PPTs in IoT architectures, followed by the edge and then fog. However, many papers did not specify which computing layer they deployed PPT in, as shown in Figure 10.
Figure 10: Publications on PPTs implementation at different computing layers in IoT architecture from 2010 to 2025
8.5 Privacy Types provided by PPTs in IoT systems
As gathered in this study, the three main privacy types focused on in most papers are user privacy, device privacy, and location privacy. User data privacy protects sensitive information, such as Personally Identifiable Information (PII), Protected Health Information (PHI), or any other data that can reveal a user’s identity. The protection of data stored, processed, or transmitted across devices that could reveal either user or device identity in IoT is called device data privacy. For example, most publications in IIoT primarily focus on the collection, processing, and transmission of device data and data related to industry personnel. Application of PPT to focus on preserving the identity of these production devices could protect not only the device’s identity but also its known vulnerabilities, which malicious threat actors could exploit. Lastly, location privacy protects IoT users against privacy threats, such as localization and tracking. This threat also applies to mobile IoT devices, including wearable and stationary devices.
Apart from these three privacy types commonly found in IoT systems, some studies have proposed PPTs that simultaneously mitigate both user and device privacy. Meanwhile, some studies have presented the general privacy of the entire IoT system without specifying any particular type of privacy.
Our observations: This study presents different privacy types proposed from the literature on PPT implementation in IoT systems. User data privacy has the highest number of papers focusing on user data protection. Few papers proposed PPT to protect device data and location privacy, as shown in Table 5 and Figure 11. This answers our research question RQ6.
Figure 11: Papers on Privacy protection types in IoT systems from 2010 to 2025
We observed that device data privacy, as proposed by most studies, still focuses on user data being stored, processed, or transmitted across IoT devices. However, any successful threat to the device’s identity privacy could compromise both the user’s and the device’s data stored in the device. Therefore, it is essential to protect device identity and sensitive device data, such as IP addresses, MAC addresses, operating systems, and serial numbers, among others, that could reveal device identity. These sensitive data associated with IoT devices could be linked with other data to identify these devices in the IoT system.
9 Step 5: Collating, Summarizing, and Reporting Results that Address Research Questions
Using the PRISMA-ScR guidelines for selecting relevant papers, our analysis reveals a steady increase in publications on PPTs from 2013 to 2019, with 2019 having the highest number of publications, as shown in Figure 8. There was a noticeable decline in 2020, probably due to the global pandemic. However, publication rates rebounded and continued to rise from 2021 to the end of 2025, when the paper collection was concluded. The following subsections present results that address the research questions outlined in section 5.
9.1 RQ1: What technologies are prominently used to develop PETs in IoT systems?
To address RQ1, this study analyzed seven technologies used to build privacy-enhancing technologies (PETs), including blockchain, Cryptography, Machine Learning and Deep Learning, Mathematical Models, Optimization, Infrastructure-Based techniques, and post-quantum algorithms. It has also been revealed that cryptography is the most prevalent technology used for privacy preservation in IoT applications, followed closely by blockchain. As observed in this study, the two publications on post-quantum algorithms are primarily applied to Industrial IoT (IIoT) and e-commerce IoT, respectively, indicating their pioneering role and potential for future research in IIoT privacy. However, informed consent management is the only non-technical approach to preserving privacy, as our study suggests.
9.2 RQ2: What are the dominant PPT’s privacy goals in IoT applications?
Our study’s research question (RQ2) examines the primary objectives of PPTs across various IoT applications. Our analysis identifies various PPT goals, including anonymity, data flow privacy, access control, privacy awareness, data isolation, and data minimization, as shown in Table 2. Among these goals, anonymity, particularly through methods such as data masking, data perturbation, and pseudonymization, emerges as the most prevalent objective in the literature, with a significant percentage of publications focusing on these PPT approaches.
9.3 RQ3: What are the commonly known and recent IoT applications added to the literature?
IoT applications commonly span cities, transportation, health, and agriculture. However, new application areas like Social IoT(S-IoT) Internet Multimedia of Things (IoMT) and Artificial Intelligence of Things (AIoT) are gaining traction as shown in Figure 12. These new platforms directly interact with users and could be vulnerable to privacy threats due to their extensive internet exposure.
Table 4: Papers on PPTs and Layer of Implementation in IoT architecture from 2010 to 2025.
IoT Application Edge Fog Cloud Local Edge & Cloud Fog & Cloud Fog, Edge & Cloud Local, Edge, Fog & Cloud Local & Cloud Not Mentioned
Smart Home [628], [397], [416], [518], [175], [92], [321], [385] [44], [382] [372], [500], [460], [461], [565], [381] [158], [104], [9] [384] [325] [232], [374], [598], [508], [354], [371], [383], [406], [624], [540], [136]
Industrial IoT [549], [186], [360], [194] [272], [188] [121], [267], [192] [287], [480], [198], [523] [122], [415] [526] [203] [434], [185], [289], [582], [140], [403], [367], [405], [211], [429], [148], [257], [314], [251]
Smart Grid [97], [77], [357], [638] [607] [470], [340], [441], [425] [103], [541], [268] [529], [273], [455], [14] [463] [512], [125], [591], [597], [88], [255], [114], [89], [281], [620], [230], [609], [349], [258]
Medical IoT [511], [491], [131], [210], [225], [233], [351], [252], [82], [49], [128], [342] [144], [417], [241], [335] [532], [472], [183], [604], [629], [496], [10], [566], [201], [622], [329], [615], [305], [172], [151], [446], [535], [471], [307], [288], [179], [220], [216], [43], [394], [20] [119], [450], [539], [260] [418] [490] [284], [424], [181] [214], [187], [332], [612], [58], [239], [242], [234], [129], [396], [442], [528], [333], [505], [451], [609], [98], [379], [73], [116], [428], [50], [102], [248], [177], [286], [147], [380], [444], [516], [68]
Social IoT [449], [317], [318], [466] [21], [611] [423], [422], [205], [134], [117]
IoV [569], [8], [590], [327] [356], [348], [553], [29], [141] [93], [120], [72], [202], [41], [94], [644], [294], [585], [271], [18] [153], [625], [291], [458], [537], [584], [65] [69], [391], [473], [521], [457], [256] [235] [498], [40], [83], [197] [246], [402], [480], [38], [13], [635], [32], [51], [319], [250], [182], [616], [515], [482], [401], [571], [259], [555], [483], [227], [536], [347], [636], [443], [567], [277], [309], [228], [452], [261], [419], [337], [495], [96], [282], [326], [157], [369], [295], [477], [189], [456], [48], [229], [527], [528], [322], [137]
Wearable IoT [407] [223], [601], [31], [331], [76] [84], [320] [623], [79], [123], [547]
Smart City [557], [169], [398], [513], [479], [168], [519] [575], [5], [2], [64] [170], [173], [334], [627] [171] [285], [154], [344], [546] [453], [30], [489], [118], [39], [399], [159], [411], [274], [215], [468], [497], [127]
IoMT [278] [145] [162]
Agriculture [578] [445], [632], [316]
Education [556]
E-commerce [200] [100], [594]
General IoT [551], [435], [204], [634], [150], [111], [550], [387], [579], [231], [637], [614], [503], [542], [522], [153], [546], [149], [237], [572], [573], [440], [586], [643], [617], [152], [454], [528], [574], [75], [292] [432], [599], [304], [80], [313], [19], [439], [438], [393], [217], [56] [167], [221], [346], [408], [595], [580], [23], [63], [531], [619], [641], [577], [306], [253], [389], [462], [46], [589], [613], [90], [524], [279], [28], [174], [263] [199], [626], [86], [373], [206], [501], [105], [67], [534], [352], [243], [339] [543], [247], [61], [27], [618], [554], [53], [538], [552] [308] [106] [87], [605], [160], [238] [392] [563], [283], [570], [264], [506], [300], [583], [530], [310], [85], [370], [427], [390], [465], [576], [324], [485], [639], [562], [298], [412], [520], [47], [378], [17], [358], [436], [433], [364], [386], [196], [561], [587], [484], [270], [275], [302], [218], [486], [7], [560], [643], [12], [549], [11], [290], [130], [608], [488], [517], [315], [525], [559], [276], [55], [336], [165], [476], [146], [404], [474], [303], [224], [596], [312], [376], [101], [240], [42], [236], [564], [593], [481], [296], [3], [464], [388], [222], [81], [630], [262], [618], [265], [603], [293], [588], [126], [592], [341]
9.4 RQ4: What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT been most implemented in the literature?
IoT applications have increased recently, with medical IoT having the highest PPTs implementation as seen in Figure 12. This shows the importance of privacy in medical IoT systems. This level of importance is a result of the sensitive medical data transactions involved in the MIoT system, which contains data such as Personally Identifiable Information (PII) and Protected Health Information (PHI). In the event of a data breach, this sensitive medical data can reveal an individual’s identity in relation to their medical condition. The next prominent IoT application is the IoV, which has a greater number of publications on PPT in recent years, as seen in Figure 12. This growth may result from the rise in the number of electric vehicles and driverless cars used in shared rides. Some of these vehicles could reveal the locations of their passengers or take the wrong route in the event of a privacy breach that disrupts their operations. Furthermore, our observation shows that PPTs are more implemented around the cloud layer in the IoT architecture in most IoT applications compared to other computing layers, as shown in Table LABEL:table:IoT-Architecturetable.
@@ 9.3 RQ3: What are the commonly known and recent IoT applications added to the literature?
Figure 12: Papers on PPT implementation in IoT applications and Technology used to build PETs from 2010 to 2025
Table 5: Publications on privacy protection types in IoT systems from 2010 to 2025.
@@
User Data Privacy 332 [170, 125, 570, 264, 557, 595, 628, 340, 591, 489, 41, 392, 232, 402, 144, 599, 480, 580, 304, 23, 460, 397, 223, 80, 612, 629, 86, 530, 310, 374, 85, 531, 145, 550, 390, 297, 465, 496, 461, 119, 641, 231, 324, 565, 417, 334, 308, 100, 639, 94, 575, 38, 298, 412, 329, 13, 635, 247, 598, 384, 32, 556, 353, 508, 491, 305, 358, 172, 10, 566, 601, 436, 354, 644, 433, 194, 294, 589, 364, 386, 201, 613, 187, 423, 289, 348, 542, 551, 141, 30, 218, 241, 150, 162, 422, 561, 529, 578, 549, 587, 446, 79, 302, 90, 332, 373, 319, 622, 203, 346, 408, 204, 87, 634, 150, 283, 372, 272, 203, 122, 470, 532, 472, 183, 604, 214, 535, 449, 69, 93, 169, 432, 221, 563, 167, 486, 416, 381, 77, 92, 522, 546, 149, 518, 31, 131, 210, 418, 5, 391, 611, 382, 473, 237, 501, 471, 572, 573, 560, 233, 175, 586, 643, 351, 524, 307, 58, 12, 288, 597, 252, 273, 549, 239, 616, 11, 290, 105, 457, 88, 482, 571, 331, 179, 605, 234, 517, 259, 129, 538, 534, 450, 396, 220, 344, 284, 216, 383, 399, 160, 357, 255, 82, 490, 585, 559, 539, 528, 55, 43, 411, 407, 483, 321, 152, 274, 260, 333, 505, 451, 455, 367, 227, 335, 106, 636, 303, 316, 49, 224, 443, 596, 376, 609, 98, 379, 73, 116, 101, 104, 454, 607, 394, 428, 89, 320, 76, 123, 528, 627, 261, 317, 466, 325, 463, 458, 9, 352, 3, 50, 102, 248, 179, 238, 620, 128, 230, 75, 519, 590, 624, 385, 618, 547, 603, 293, 349, 177, 327, 258, 425, 588, 126, 424, 20, 136, 268, 342, 638, 286, 181, 609, 147, 380, 444, 117, 65, 516, 68]
Device Data Privacy 45 [626, 506, 543, 370, 313, 485, 47, 196, 435, 199, 371, 287, 185, 188, 67, 415, 401, 140, 186, 28, 403, 192, 165, 211, 114, 429, 291, 148, 452, 314, 360, 251, 198, 523, 497, 96, 477, 526, 137, 341]
User and Device Data Privacy 51 [453, 285, 387, 623, 619, 389, 462, 520, 19, 503, 439, 438, 44, 484, 270, 275, 500, 434, 121, 512, 511, 21, 225, 441, 554, 8, 242, 40, 555, 174, 158, 536, 347, 256, 14, 197, 228, 406, 215, 263, 419, 337, 194, 540, 127, 552, 369, 295, 229, 528]
Location Privacy 45 [300, 202, 583, 576, 577, 29, 615, 17, 614, 84, 278, 562, 51, 205, 246, 553, 153, 206, 182, 118, 618, 515, 525, 336, 146, 271, 567, 277, 309, 134, 64, 388, 495, 282, 157, 189, 456, 48, 527, 118, 322]
General Privacy 99 [111, 427, 579, 97, 306, 253, 398, 61, 46, 378, 637, 267, 569, 356, 120, 72, 173, 171, 582, 625, 250, 625, 498, 27, 521, 130, 56, 513, 594, 279, 217, 608, 53, 488, 315, 39, 546, 445, 617, 507, 276, 159, 476, 632, 83, 503, 103, 541, 404, 405, 474, 2, 312, 240, 42, 170, 236, 257, 564, 593, 200, 468, 318, 18, 481, 296, 574, 243, 464, 281, 222, 479, 168, 339, 81, 630, 609, 262, 265, 326, 292, 537, 584, 592]
@@
9.5 RQ5: What privacy types are the main focus of most PPTs in IoT systems?